Martinez Company accumulates the following data concerning a mixed cost, using units produced as the activity level. \Martinez Company accumulates the following data concerning a mixed cost, using units produced as the activity level. March Ap

Show Work

1. Compute the variable cost per unit using the high-low method. (Round variable cost per unit to 2 decimal places e.g. 12.25.)
2. Compute the fixed cost elements using the high-low method. (Round answer to 0 decimal places, e.g. 5,275.)

3. Estimate the total cost if the company produces 8,100 units. (Round answer to 0 decimal places, e.g. 5,275.)
Answer #1

1. Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (28,050-24,042)/(10,600-8930)

= 2.4

2. Fixed cost = Total cost - Variable cost

= 28,050 - (2.4*10,600)

= 2610

3. Total Cost

= (2.4*8100)+2610

= 22,050